Swifty Products, a rapidly growing distributor of home gardening equipment, is formulating its plans for the coming year. Carol Jones, the firm’s marketing director, has completed the following sales forecast.

Month Sales Month Sales
January $904,900 July $1,501,400
February $1,002,000 August $1,501,400
March $904,900 September $1,606,000
April $1,154,600 October $1,606,000
May $1,260,000 November $1,501,400
June $1,406,000 December $1,704,000


Phillip Smith, an accountant in the Planning and Budgeting Department, is responsible for preparing the cash flow projection. He has gathered the following information.

All sales are made on credit.
Swifty’s excellent record in accounts receivable collection is expected to continue, with 60% of billings collected in the month after sale and the remaining 40% collected two months after the sale.
Cost of goods sold, Swifty’s largest expense, is estimated to equal 40% of sales dollars. Seventy percent of inventory is purchased one month prior to sale and 30% during the month of sale. For example, in April, 30% of April cost of goods sold is purchased and 70% of May cost of goods sold is purchased.
All purchases are made on account. Historically, 75% of accounts payable have been paid during the month of purchase, and the remaining 25% in the month following purchase.
Hourly wages and fringe benefits, estimated at 30% of the current month’s sales, are paid in the month incurred.
General and administrative expenses are projected to be $1,575,000 for the year. A breakdown of the expenses follows. All expenditures are paid monthly throughout the year, with the exception of property taxes, which are paid in four equal installments at the end of each quarter.
   Salaries and fringe benefits $ 323,700
   Advertising 378,800
   Property taxes 142,400
   Insurance 195,600
   Utilities 184,800
   Depreciation 349,700
   Total $ 1,575,000
Operating income for the first quarter of the coming year is projected to be $329,100. Swifty is subject to a 40% tax rate. The company pays 100% of its estimated taxes in the month following the end of each quarter.

Swifty maintains a minimum cash balance of $50,000. If the cash balance is less than $50,000 at the end of the month, the company borrows against its 12% line of credit in order to maintain the balance. All borrowings are made at the beginning of the month, and all repayments are made at the end of the month (in increments of $1,000). Accrued interest is paid in full with each principal repayment. The projected cash balance on April 1 is $51,300

Prepare the cash payments budget for the second quarter. (Round answers to 0 decimal places, e.g. 5,275. Enter answers in necessary fields only. Leave other fields blank. Do not enter 0.)

Cash Payments Budget
April May June
March purchases $enter a dollar amountenter a dollar amount

$enter a dollar amountenter a dollar amount

$enter a dollar amountenter a dollar amount

April purchases enter a dollar amountenter a dollar amount

enter a dollar amountenter a dollar amount

enter a dollar amountenter a dollar amount

May purchases enter a dollar amountenter a dollar amount

enter a dollar amountenter a dollar amount

enter a dollar amountenter a dollar amount

June purchases enter a dollar amountenter a dollar amount

enter a dollar amountenter a dollar amount

enter a dollar amountenter a dollar amount

$Enter a total amountEnter a total amount

$Enter a total amountEnter a total amount

$Enter a total amountEnter a total amount

Accounts Payable balance at the end of second quarter of 2015 $Enter a dollar amount of the Accounts Payable balance at the end of second quarter of 2015Enter a dollar amount of the Accounts Payable balance at the end of second quarter of 2015
Answer #1
Cash payments budget for Purchases
Calculation
April May June Quarter April May June
Mar Purchases       107,969       107,969 =((904,900*40%*30%)+(1,154,600*40%*70%))*25%
Apr Purchases       368,514       122,838       491,352 =491,352*75% =491,352*25%
May Purchases       408,660       136,220       544,880 =544,880*75% =544,880*25%
June Purchases       441,834       441,834 =589,112*75%
Total disbursements for purchases       476,483       531,498       578,054    1,586,035
Accounts Payable balance at end of 2nd quarter of 2015       147,278 =589,112*25%

Note: First prepare below Inventory Purchase Budget for Purchases figures, then make above Cash payments budget !

Inventory Purchase budget Calculation
April May June Quarter April May June
April COGS       138,552       138,552 =1,154,600*40%*30%
May COGS       352,800       151,200       504,000 =1,260,000*40%*70% =1,260,000*40%*30%
June COGS       393,680       168,720       562,400 =1,406,000*40%*70% =1,406,000*40%*30%
July COGS       420,392       420,392 =1,501,400*40%*70%
Required Purchases       491,352       544,880       589,112    1,625,344
Cash receipts Calculation
April May June Quarter April May June
Feb sales       400,800       400,800 =1,002,000*40%
Mar sales       542,940       361,960       904,900 =904,900*60% =904,900*40%
Apr sales       692,760       461,840    1,154,600 =1,154,600*60% =1,154,600*40%
May sales       756,000       756,000 =1,260,000*60%
Total collections from sales       943,740    1,054,720    1,217,840    3,216,300
Accounts Receivable balance at end of 2nd quarter of 2015    1,910,000 =(1,260,000*40%)+1,406,000
Cash budget
April May June Quarter April May June
Beginning Cash balance          51,300          50,295          50,196          51,300
Add: Collections from sales       943,740    1,054,720    1,217,840    3,216,300
Total cash available to spend       995,040    1,105,015    1,268,036    3,267,600
Less: Cash payments                   -  
Inventory purchases       476,483       531,498       578,054    1,586,035
Wages and fringe benefits       346,380       378,000       421,800    1,146,180 =1,154,600*30% =1,260,000*30% =1,406,000*30%
Salaries and fringe benefits          26,975          26,975          26,975          80,925 =323,700/12 =323,700/12 =323,700/12
Advertising          31,567          31,567          31,567          94,700 =378,800/12 =378,800/12 =378,800/12
Property taxes                   -                     -            35,600          35,600 =142,400/4
Insurance          16,300          16,300          16,300          48,900 =195,600/12 =195,600/12 =195,600/12
Utilities          15,400          15,400          15,400          46,200 =184,800/12 =184,800/12 =184,800/12
Income taxes       131,640                   -                     -         131,640 =329,100*40%
Total cash payments    1,044,745       999,740    1,125,696    3,170,180
Cash Excess (deficiency)       (49,705)       105,276       142,340          97,420
Minimum cash balance          50,000          50,000          50,000          50,000
Cash Excess (needed)       (99,705)          55,276          92,340          47,420 -41,247-50,000 =112,225-50,000 =141,903-50,000
Financing
Borrowings at beginning       100,000       100,000 Rounded off 99,705 to 100,000 for borrowings calculation
Repayments at end       (54,000)       (46,000)     (100,000)
Interest @ 12%          (1,080)          (1,380)          (2,460) =54,000*12%*2/12 =46,000*12%*3/12
Total Financing       100,000       (55,080)       (47,380)          (2,460)
Ending Cash balance          50,295          50,196          94,960          94,960 -49,705+100,000 =105,276-55,080 =142,340-47,380
Cash Excess (deficiency) + Total Financing